\int _1^2\int _2^3\int _0^1\:8xyzdzdxdy


1
Expert's answer
2021-02-24T06:07:41-0500

"\displaystyle\int\limits _1^2\int\limits _2^3\int\limits _0^1\:8xyz\:dzdxdy="


"\displaystyle8\int\limits _1^2\int\limits _2^3\int\limits _0^1\:xyz\:dzdxdy="


"\displaystyle8\int\limits _1^2\int\limits _2^3\left[\int\limits _0^1\:xyz\:dz\right]dxdy="


"\displaystyle8\int\limits _1^2\int\limits _2^3\left[xy\int\limits _0^1\:z\:dz\right]dxdy="


"\displaystyle8\int\limits _1^2\int\limits _2^3xy\left[\int\limits _0^1\:z\:dz\right]dxdy="


"\displaystyle8\int\limits _1^2\int\limits _2^3xy\left[\frac{z^2}{2}\right]_0^1dxdy="


"\displaystyle8\int\limits _1^2\int\limits _2^3xy\left[\frac{1^2}{2}-\frac{0^2}{2}\right]dxdy="


"\displaystyle8\int\limits _1^2\int\limits _2^3xy\cdot\frac{1}{2}\:dxdy="


"\displaystyle8\cdot\frac{1}{2}\int\limits _1^2\int\limits _2^3xy\:dxdy="


"\displaystyle4\int\limits _1^2\left[\int\limits _2^3xy\:dx\right]dy="


"\displaystyle4\int\limits _1^2\left[y\int\limits _2^3x\:dx\right]dy="


"\displaystyle4\int\limits _1^2y\left[\int\limits _2^3x\:dx\right]dy="


"\displaystyle4\int\limits _1^2y\left[\frac{x^2}{2}\right]_2^3\:dy="


"\displaystyle4\int\limits _1^2y\left[\frac{3^2}{2}-\frac{2^2}{2}\right]\:dy="


"\displaystyle4\int\limits _1^2y\left[\frac{9-4}{2}\right]\:dy="


"\displaystyle4\int\limits _1^2y\cdot\frac{5}{2}\:dy="


"\displaystyle4\cdot\frac{5}{2}\int\limits _1^2y\:dy="


"\displaystyle10\int\limits _1^2y\:dy="


"\displaystyle10\left[\frac{y^2}{2}\right]_1^2="


"\displaystyle10\cdot\left[\frac{2^2}{2}-\frac{1^2}{2}\right]="


"\displaystyle10\cdot\frac{4-1}{2}="


"\displaystyle10\cdot\frac{3}{2}=" "15"


Answer:

"\displaystyle\int\limits _1^2\int\limits _2^3\int\limits _0^1\:8xyz\:dzdxdy=15"